how do you solve this problem (x)/(4)+5=4+(x)/(2)
Answer by nerdybill(7384)

(x)/(4)+5=4+(x)/(2)
x/4+5=4+x/2
multiply both sides by 4:
x+20=16+2x
20=16+x
4 = x